# On the Feasibility of Simple Transformer for Dynamic Graph Modeling

### 25 Jan 2024 v1

@@ 4.1. Temporal Ego-graph
As mentioned earlier, the sequence modeling capability of the Transformer architecture is well-suited for dynamic graph modeling. The strategy of mapping dynamic graphs into a sequence of tokens is crucial for the supported features and computational complexity. In this paper, we regard nodes in the dynamic graphs as input tokens which is a common approach in Transformer models for graphs. Besides, to preserve more historical interactions of all the nodes and ensure the scalability of dealing with large receptive fields, we extract the temporal ego-graph for each node in the dynamic graph. These temporal graphs are mapped into sequences to capture the structural and temporal evolution.
Specifically, we denote as an ego-node in the temporal graph . We extract the historically interacted nodes for and concatenate them into a sequence as input for Transformer architecture. Formally, we denote the temporal ego-graph for the ego-node as , where is the length of the historical interactions for node . To better model the patterns within the input sequence, we follow similar practices as in NLP sequence modeling tasks and include some special tokens designed for our task. Finally, the input sequence and output sequence are constructed as follows 222Special tokens in the beginning and at the end such as “” are omitted for easy illustration.:
where the “” and “” are special tokens indicating the start and end of the input historical sequence. The “” and “” are reserved for predicting the next nodes at a future time. Specifically, the model will halt its predictions once the end special token is generated, enabling automatic decisions on the number of future interactions.
@@ 4.2. Temporal Alignment
In the original Transformer architecture, the input sequence is treated as a sequence of tokens, and the model captures the relationships between these tokens based on their relative positions in the sequence, representing temporal order information. However, it inherently lacks the capability to account for the universal time domain and the time interval and frequency information. In pursuit of this objective, we segment the time domain into discrete, coarse-grained time steps, with each time step representing the same time interval, such as one week or one month, determined by dataset characteristics. It’s important to note that our approach differs from discrete-time graph modeling, as within each time step, we consider the precise temporal order of each link. We introduce a straightforward yet effective strategy to incorporate temporal alignment within dynamic graphs into the input sequence of the Transformer architecture. This strategy entails the use of special temporal tokens explicitly denoting different time steps that are globally recognized across all nodes. Suppose we split the time domain into time steps, the sequence of ego-node in time step is denoted as follows:
where represents the historical sequence of node at time step whith length of . are temporal tokens that serve as indicators of temporal alignment, allowing the model to recognize and capture temporal patterns in the data. By doing so, our approach enhances the Transformer’s ability to understand the dynamics of the dynamic graph, making it more effective in tasks like predicting future interactions in social networks or other dynamic systems where temporal patterns are crucial.
@@ 4.3. Training objective
A training sample is formed by concatenating the input and output as . We denote it as with tokens. For a given training instance in this format, we follow the original masking strategy, where, during the prediction of the -th token, only the input sequence up to position is taken into account, while the subsequent tokens are subject to masking. The joint probability of the next token is calculated as follows:
where is the generated sequence before step . denotes the probability distribution of the token to be predicted at step conditioned with the tokens . It is computed as:
where LN means layer normalization. denotes the hidden representation of the historically generated tokens before step , which is obtained by the last layer of Transformer. is the learned parameter aiming to compute the probability distribution across the vocabulary of nodes in the graph.
Given a dataset containing training instances, the loss function for training the model with parameters is defined as the negative log-likelihood over the entire training dataset as follows:
where is the length of the instance .
We outline the training procedure of SimpleDyG in Algorithm 1. For each prediction step of one training instance, the hidden representations of the generated sequence are used for predicting the next token. The joint probability of the next token is computed using Equations 6 and 7. Our model is trained using the Adam optimizer with a loss function based on negative log-likelihood, as presented in Equation 8.
@@ 5.1. Experimental Setup
Datasets. To evaluate the performance of our proposed method, we conducted experiments on four datasets from various domains, including the communication network UCI (Panzarasa et al., 2009), the rating network ML-10M (Harper and Konstan, 2015), the citation network Hepth (Leskovec et al., 2005), and the multi-turn conversation dataset MMConv (Liao et al., 2021). The detailed statistics of all datasets after preprocessing are presented in Table 1.
@@
UCI (Panzarasa et al., 2009): it represents a social network in which links represent messages exchanged among users. For temporal alignment, we and divide the dataset into 13 time steps following (Sankar et al., 2020).
ML-10M (Harper and Konstan, 2015): we utilized the ML-10M dataset from MovieLens dataset comprising user-tag interactions, where the links connect users to the tags they have assigned to specific movies. For temporal alignment, the dataset is split into 13 time steps following (Sankar et al., 2020).
Hepth (Leskovec et al., 2005): it is a citation network related to high-energy physics theory. We extract 24 months of data from this dataset and split them into 12 time steps for temporal alignment. Note that this dataset contains new emerging nodes as time goes on. We use the extra word2vec (Mikolov et al., 2013) model to extract the raw feature for each paper based on the abstract.
MMConv (Liao et al., 2021): this dataset contains a multi-turn task-oriented dialogue system that assists users in discovering places of interest across five domains. Leveraging this rich annotation, we represent the dialogue as a dynamic graph which is also a widely studied approach in task-oriented dialogue systems. For temporal alignment, we empirically divided the dataset into 16 time steps, each corresponding to a distinct turn in the conversation.
@@
Baselines. We compare our method with baselines in two categories: (1) discrete-time approaches: DySAT (Sankar et al., 2020) and EvolveGCN (Pareja et al., 2020) (2) continuous-time approaches: DyRep (Trivedi et al., 2019), JODIE (Kumar et al., 2019), TGAT (Xu et al., 2020), TGN (Rossi et al., 2020), TREND (Wen and Fang, 2022) and GraphMixer (Cong et al., 2022).
@@
EvolveGCN (Pareja et al., 2020) adapts to evolving graph structures by employing RNN to evolve graph convolutional network parameters.
@@
JODIE (Kumar et al., 2019) focuses on modeling the binary interaction among users/items by two coupled RNNs. A projection operator is designed to predict the future representation of a node at any time.
@@
TGN (Rossi et al., 2020) combines the memory modules and message-passing to maintain the dynamic representations. This model also adopts time encoding and temporal graph attention layers.
TREND (Wen and Fang, 2022) exploits Hawkes process-based GNN for temporal graph representation learning. It integrates event and node dynamics to fully capture the individual and collective characteristics of events.
GraphMixer (Cong et al., 2022) relies on MLP layers and neighbor mean-pooling to learn the link and node encoders. An offline time encoding function is adopted to capture the temporal information.
Implementation Details. In this paper, we evaluate the performance of SimpleDyG on the link prediction task. Given the ego-nodes, the objective of the link prediction task is to predict the possible linked nodes at time step . For all the datasets, we follow the setting in (Cong et al., 2022) by treating the temporal graphs as undirected graphs. We split each dataset into training/validation/testing based on the predefined time steps. We choose the data at the last time step as the testing set, while the data at time step serves as the validation set, with the remaining data for training. We tune the parameters for all methods on the validation set. All experiments are repeated ten times, and we report the averaged results with standard deviation. We provide further implementation details and hyper-parameter settings for the baselines in Appendix A and B.
Evaluation Metrics. In our evaluation, we carefully selected metrics that are well-suited to our specific task. The goal of the link prediction task is to predict a set of nodes linked to each ego-node. Notably, our SimpleDyG model predicts a node sequence, with each prediction influenced by the prior ones until the generation of an end token. In contrast, the baseline models make simultaneous predictions of entire ranking sequences for each ego-node. To evaluate ranking performance and set similarity between predicted and ground truth node sets, we employ two key metrics: NDCG@5 and Jaccard similarity. NDCG@5 is a well-established metric commonly used in information retrieval and ranking tasks (Wang et al., 2019a), aligning with our objective of ranking nodes and predicting the top nodes linked to an ego-node. On the other hand, Jaccard similarity is valuable for quantifying the degree of overlap between two sets (Jaccard, 1901), measuring the similarity between predicted nodes and the ground truth nodes associated with the ego-node. Specifically, for the baseline models, we choose the top k nodes () as the predicted set, as they are not generation models and cannot determine the end of the prediction. We then select the maximum Jaccard similarity value across different k as the final Jaccard similarity score. This comprehensive evaluation strategy ensures a thorough assessment of our models and baselines in predicting linked nodes.
@@
5.2. Main Performance Comparison
@@
Firstly, we find that continuous-time approaches generally perform better than discrete ones across a wide range of scenarios, indicating the important role of time-related information in dynamic graph analysis. Notably, continuous-time baselines such as GraphMixer exhibit superior performance. This superiority can be mainly attributed to the simple MLP-Mixer architecture, which makes it easier to capture long-term historical sequences with lower complexity. In contrast, other models like DyRep, TGAT, and TGN, which rely on complex designs such as GNNs and GATs, display subpar performance. This phenomenon stems from the inherent limitations of GNNs and GATs in capturing distant relationships or broader historical contexts within predefined time windows.
Secondly, for the inductive scenarios such as the Hepth dataset, the models deployed by GNNs, GATs, and Transformer show advanced performance. Their effectiveness lies in their ability to capture intricate patterns and relationships within dynamic graphs, especially when faced with inductive scenarios. In contrast, simpler models may struggle to adapt to these situations, resulting in suboptimal performance.
In summary, while continuous-time approaches have generally been shown effective, it is essential to consider the specific characteristics of the application scenario and strike a delicate balance between model complexity and the necessity to capture long-range dependencies in dynamic graph modeling.
@@ 5.3. Effect of Extra Tokens
We design extra tokens to make the vanilla Transformer architecture more suitable for dynamic graph modeling. To assess their effectiveness, we conduct an in-depth analysis of these token designs including special tokens indicating the input and output, and the temporal tokens for aligning among temporal ego-graphs.
@@
Impact of the special tokens. The special tokens include the start and end of the historical sequence (“” and “”), as well as the predicted sequence (“”and “”). To comprehensively evaluate their effect across diverse scenarios, we examine two degenerate variants: (1) same special, where we use the same special tokens for input and output. (2) no special, where we entirely removed all special tokens from each sample. We show the results in Table 3 and make the following observations.
@@
In general, special tokens enhance the link prediction performance across different datasets. Furthermore, the differences between the same special and original SimpleDyG tend to be minimal. However, an interesting finding emerges in the case of the Hepth dataset, where the no special scenario yields the best performance. It can be explained by the specific character of the citation dataset. In the testing data of Hepth, the ego-nodes are all newly emerged nodes indicating the newly published papers. Consequently, the input samples lack any historical information, leaving the distinction between history and the future meaningless.
@@
Impact of temporal tokens. To comprehensively evaluate the impact of temporal tokens, we compare the performance with two degenerate variants: (1) same time, where we do not distinguish specific time steps and employ the same temporal tokens for each time step. (2) no time, in which we entirely removed all temporal tokens from each sample. The results are presented in Table 4 and we have the following observations.
@@
It is surprising and interesting to observe performance improvement with a simpler design for temporal alignment. This phenomenon is most noticeable in the MMConv and Hepth datasets due to the characteristics of these datasets. The citation relationship and the conversation among different ego-nodes do not strictly follow temporal segmentation. Using the same temporal tokens or none at all allows the model to adapt more naturally to this temporal order.The temporal alignment plays an important role for UCI and ML-10M datasets. However, they show different trends with the same time version. The reason is that the communication habits of different users are more related to temporal information, while the rating habits of users are more subjective in ML-10M dataset.
5.4. The Performance of Multi-step Prediction
@@
We evaluate the ability of SimpleDyG for multi-step prediction with the time steps range from to , utilizing a model that has been trained on data up to time . Here, the time step means the coarse segment of the time domain as did in the temporal alignment. In our experiment, we set as three and achieve multi-step prediction step by step constrained by the results of previous steps. The performance trends of SimpleDyG with two baselines TGAT and GraphMixer are illustrated in Figure 3.
We observe a natural decay in performance over time for all methods, as anticipated. However, what stands out is SimpleDyG’s ability to consistently outperform the baselines as time progresses. This observed trend underscores the effectiveness of our proposed Transformer architecture in modeling dynamic graph data. Notably, different datasets exhibit varying patterns of performance decay over time, highlighting the importance of dataset-specific considerations in dynamic graph analysis. For instance, in the case of the ML-10M and Hepth datasets, we notice a relatively slight performance drop over time. This phenomenon can be attributed to specific characteristics inherent to these datasets. In the ML-10M dataset, the presence of numerous historical interactions contributes to a relatively stable performance trend. The dataset’s richness in historical data allows the model to absorb small noise or fluctuations without a significant impact on overall performance. On the other hand, the Hepth dataset introduces a unique challenge due to the presence of new nodes at each time step. Despite this inherent complexity, SimpleDyG still demonstrates its adaptability by maintaining competitive performance, reflecting its capability to adapt to dynamic scenarios and evolving graph structures.

### 27 Feb 2024 v2

@@ 4.1. Temporal Ego-graph
The strategy of mapping a dynamic graph into a sequence of tokens is crucial for utilizing the Transformer architecture for dynamic graph modeling. In this work, we regard nodes in the dynamic graphs as input tokens, which is a common approach in Transformer models for graphs. Besides, to preserve the historical interactions of all the nodes while ensuring scalability, we extract a temporal ego-graph for each node in the dynamic graph. Each ego-graph serves as the receptive field of its ego-node, which is mapped into a sequence to capture the structural and temporal evolution of the ego-node.
Specifically, we denote as an ego-node in the dynamic graph . We extract the historically interacted nodes for and construct an temporal ego-graph centered at . Formally, we denote the temporal ego-graph for the ego-node as a chronologically ordered sequence , where is the length of the sequence. Note that , and represent some historical interactions and , respectively, such that .
To better model the patterns within the sequences, we follow similar practices as in natural language processing and include some special tokens designed for our task. During training, the input sequence and the ground-truth output sequence are constructed as follows111Special tokens in the beginning and at the end such as “” are omitted for easy illustration..
where the “” and “” are special tokens indicating the start and end of the input historical sequence, and “” and “” are reserved for signalling the prediction of the next nodes following the input sequence. Specifically, the model will halt its predictions once the end special token is generated, enabling automatic decisions on the number of predictions at a future time point. Potentially, by devising appropriate special tokens, our framework can also support other operations such as link deletion, as illustrated in Appendix A.
@@ 4.2. Temporal Alignment
In the original Transformer, the input sequence is treated as a sequence of tokens, and the model captures the relationships between these tokens based on their relative positions in the sequence. For dynamic graph modeling, the positions in the sequence represent the temporal order. However, it inherently lacks the capability to account for a universal time domain across sequences, to synchronize the time interval and frequency information.
We introduce a straightforward yet effective strategy to incorporate temporal alignment into the input sequences of the Transformer. First, we segment the time domain into coarse-grained time steps, where the intervals between two consecutive time steps are equal, such as one week or one month, determined by dataset characteristics. It is important to note that our approach differs from discrete-time approaches: Within each time step, we consider the precise temporal order of each event. In contrast, discrete-time approaches treat each time step as a static snapshot. Next, we incorporate special temporal tokens into the sequences, which explicitly denote different time steps that are globally recognized across all sequences. Suppose we split the time domain into time steps, where the input sequence includes the first time steps and the output sequence covers the last time step. Then, the sequences of the ego-node can be denoted as follows:
Here represents the historical sequence of node at time step with length . In particular, “”, , “” are temporal tokens that serve as indicators of temporal alignment, enabling the model to recognize and capture temporal patterns in different sequences. The temporal tokens enhances the Transformer’s ability to understand the dynamics across the entire graph.
@@ 4.3. Training objective
A training instance is formed by concatenating the input and output as . We denote it as with tokens. For a given training instance in this format, we follow the typical masking strategy: During the prediction of the -th token, only the input sequence up to position , denoted by , is taken into account, while the subsequent tokens are subject to masking. The joint probability of the next token is calculated as follows:
where is the probability distribution of the token to be predicted at step conditioned on the tokens . It is computed as
where LN means layer normalization, denotes the hidden representation of the historically generated tokens before step , which is obtained from the last layer of the Transformer, and is a learnable matrix aiming to compute the probability distribution across the vocabulary of nodes in the graph.
Given a set of training instances , the loss function for training the model with parameters is defined as the negative log-likelihood over , as follows:
We outline the training procedure of SimpleDyG in Algorithm 1. For each prediction step of one training instance, the hidden representations of the generated sequence are used for predicting the next token. The joint probability of the next token is computed using Equations 10 and 11. Our model is trained using the Adam optimizer with a loss function based on the negative log-likelihood in Equation 12.
@@ 5. Experiments
In this section, we conduct extensive experiments on four public datasets across different domains, with comparison to the state-of-the-art baselines and detailed analysis of the model performance.
@@ 5.1. Experimental Setup
Datasets. To evaluate the performance of our proposed method, we conducted experiments on four datasets from various domains, including the communication network UCI (Panzarasa et al., 2009), the rating network ML-10M (Harper and Konstan, 2015), the citation network Hepth (Leskovec et al., 2005), and the multi-turn conversation dataset MMConv (Liao et al., 2021b). The detailed statistics of all datasets after preprocessing are presented in Table 1.
@@
UCI (Panzarasa et al., 2009): It represents a social network in which edges represent messages exchanged among users. For temporal alignment, we employ 13 time steps following previous work (Sankar et al., 2020).
ML-10M (Harper and Konstan, 2015): We utilize the ML-10M dataset from MovieLens comprising user-tag interactions, where the edges connect users to the tags they have assigned to specific movies. For temporal alignment, we employ 13 time steps following previous work (Sankar et al., 2020).
Hepth (Leskovec et al., 2005): It is a citation network for high-energy physics theory. For temporal alignment, we extract 24 months of data from this dataset and split them into 12 time steps based on the wall-clock timestamps. Note that this dataset contains new emerging nodes as time goes on. We use the word2vec model (Mikolov et al., 2013) to extract the input feature for each paper based on the abstract.
MMConv (Liao et al., 2021b): It contains a multi-turn task-oriented dialogue system that assists users in discovering places of interest across five domains. Leveraging this rich annotation, we represent the dialogue as a dynamic graph, a widely adopted strategy in task-oriented dialogue systems. For temporal alignment, we employ 16 time steps, each corresponding to a distinct turn in the conversation.
@@
Baselines. We compare SimpleDyG with baselines in two categories: (1) discrete-time approaches: DySAT (Sankar et al., 2020) and EvolveGCN (Pareja et al., 2020); (2) continuous-time approaches: DyRep (Trivedi et al., 2019), JODIE (Kumar et al., 2019), TGAT (Xu et al., 2020), TGN (Rossi et al., 2020), TREND (Wen and Fang, 2022) and GraphMixer (Cong et al., 2022).
@@
EvolveGCN (Pareja et al., 2020) employs RNNs to evolve graph convolutional network parameters.
@@
JODIE (Kumar et al., 2019) focuses on modeling the binary interaction among users and items by two coupled RNNs. A projection operator is designed to predict the future node representations at any time.
@@
TGN (Rossi et al., 2020) combines the memory modules and message passing to maintain the dynamic representations. This model also adopts time encoding and temporal graph attention layers.
TREND (Wen and Fang, 2022) exploits Hawkes process-based GNNs for dynamic graph modeling. It integrates event and node dynamics to capture the individual and collective characteristics of events.
GraphMixer (Cong et al., 2022) relies on MLP layers and neighbor mean-pooling to learn the edge and node encoders. An offline time encoding function is adopted to capture the temporal information.
Implementation Details. We evaluate the performance of SimpleDyG on the link prediction task. Given the ego-nodes, the objective of the link prediction task is to predict the possible linked nodes at time step . For all the datasets, we follow the previous setting (Cong et al., 2022) by treating the dynamic graphs as undirected graphs. We split each dataset into training/validation/testing based on the predefined time steps. We choose the data at the last time step as the testing set, while the data at time step serves as the validation set, with the remaining data for training. We tune the hyperparameters for all methods on the validation set. All experiments are repeated ten times, and we report the averaged results with standard deviation. We provide further implementation details and hyperparameter settings in Appendices B and C.
Evaluation Metrics. In our evaluation, we carefully select metrics catered to our specific task. The goal of the link prediction task is to predict a set of nodes linked to each ego-node at a given time step. Notably, our SimpleDyG model predicts a node sequence, with each prediction influenced by the prior ones until the generation of an end token. In contrast, the baseline models make simultaneous predictions of entire ranking sequences for each ego-node. To evaluate the ranking performance and the similarity between predicted and ground-truth node sets, we employ two key metrics: NDCG@5 and Jaccard similarity. NDCG@5 is a well-established metric commonly used in information retrieval and ranking tasks (Wang et al., 2019), aligning with our objective of ranking nodes and predicting the top nodes linked to an ego-node. On the other hand, Jaccard similarity is valuable for quantifying the degree of overlap between two sets (Jaccard, 1901), measuring the similarity between predicted nodes and the ground-truth nodes associated with the ego-node. Specifically, for the baseline models, we choose the top k nodes () as the predicted set, as they are not generative models and cannot determine the end of the prediction. We then select the maximum Jaccard similarity value across different k’s as the final Jaccard similarity score. This evaluation strategy ensures a thorough and fair assessment of the baselines in comparison to SimpleDyG.
@@
5.2. Performance Comparison to Baselines
@@
First, we find that continuous-time approaches generally perform better than discrete ones across a wide range of scenarios, indicating the important role of time-related information in dynamic graph analysis. Notably, continuous-time baselines such as GraphMixer exhibit superior performance. This superiority can be mainly attributed to the simple MLP-Mixer architecture, which makes it easier to capture long-term historical sequences with lower complexity. In contrast, other models like DyRep, TGAT, and TGN, which rely on complex designs such as GNNs and GATs, display subpar performance. This phenomenon stems from the inherent limitations of GNNs and GATs in capturing distant relationships or broader historical contexts within predefined time windows.
Second, in inductive scenarios like the Hepth dataset, where the ego-nodes in the testing data are newly emerged nodes, continuous-time models that employ a GNN-based backbone exhibit superior performance compared to GraphMixer. To be able to handle new nodes, the initial node features are constructed using word2vec, which might be relatively coarse. Since GraphMixer predominantly relies on an MLP-based architecture, it may encounter challenges given the coarse-grained initial features. Conversely, GNN-based methods integrate structural information with these features, thereby empowering them to excel in the inductive scenario. Nevertheless, in our Transformer-based model, there is the added advantage of modeling long-range dependencies, resulting in consistently better performance of SimpleDyG.
@@ 5.3. Effect of Extra Tokens
We design extra tokens to make the vanilla Transformer architecture more suitable for dynamic graph modeling. To assess their effectiveness, we conduct an in-depth analysis of these token designs, including special tokens that signal the input and output, and the temporal tokens that align different sequences.
@@
Impact of special tokens. The special tokens include the start and end of the historical sequence (“” and “”), as well as the predicted sequence (“”and “”). To comprehensively evaluate their effect across diverse scenarios, we examine two variants of SimpleDyG: (1) same special, where we use the same special tokens for input and output. (2) no special, where we entirely remove all special tokens from each instance. We show the results in Table 3 and make the following observations.
@@
In general, special tokens enhance the link prediction performance across different datasets. Furthermore, the differences between the same special and original SimpleDyG tend to be small. However, an interesting finding emerges in the case of the Hepth dataset, where the no special variant yields the best performance. It can be explained by the specific characteristic of the citation graph. In the testing data of Hepth, the ego-nodes are newly emerged nodes, representing the newly published papers. Consequently, the input samples lack any historical information, leaving the distinction between the history and the predicted irrelevant.
@@
Impact of temporal tokens. To evaluate the impact of temporal tokens, we compare the performance of SimpleDyG with two ablated variants: (1) same time, where we do not distinguish specific time steps and employ the same temporal tokens for each time step, and (2) no time, in which we entirely remove all temporal tokens from all sequences.
@@
The results are presented in Table 4. It is surprising and interesting to observe performance improvement with less or no temporal alignment, particularly on the MMConv and Hepth datasets. We hypothesize that the citation relationship and the conversation among different ego-nodes do not strictly follow a universal temporal framework. Using the same temporal tokens (same time) or none at all (no time) allows the model to adapt more naturally to the temporal order. The temporal alignment plays a more important role for the UCI and ML-10M datasets. However, they show different trends with the same time version. A potential reason is that, in UCI, the communication patterns between different users are sensitive to the segmentation into different time steps. Hence, same time performs the worst, as it divides a sequence into time steps yet without time differentiation to align across the sequences, where the additional same tokens may confuse the model. On the other hand, no time still retains the full temporal order, and thus perform better than same time.
5.4. Performance of Multi-step Prediction
@@
We evaluate the ability of SimpleDyG for multi-step prediction with the time steps ranging from to , utilizing a model that has been trained on data up to time . In our experiment, we set . For SimpleDyG, we perform step-by-step prediction, conditioned on the results of previous steps. We compare to two competitive baselines, TGAT and GraphMixer. As both methods incorporate time information via time encoding, we employ the model trained at time to directly predict the links at time . Their performance trends are plotted in Figure 3.
We observe a natural decay in performance over time for all methods, as anticipated. However, SimpleDyG consistently outperforms the baselines as time progresses. This trend underscores the effectiveness of our Transformer architecture in modeling long-term dependencies in a dynamic graph.